OpenStudy (anonymous):

Just need a walkthrough on what I should do first 2/3x+4=2x I think I know what I should do first but not 100% sure

OpenStudy (anonymous):

Do I multiply each side by the opposite reciprocal of 2/3?

OpenStudy (anonymous):

Are you ready?

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 }\times \frac{ 2 }{ 3 }\chi +4=2\chi \times \frac{ 3 }{2}\] Is this what I do?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Okay then what should I do first?

OpenStudy (anonymous):

Subtract 4 from both sides and tell me what you get

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }\chi =2\chi -4\]

OpenStudy (anonymous):

Now multiply both sides with 3 and tell me what you get

OpenStudy (anonymous):

\[2\chi=6\chi-12\]

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

Now subtract 6x on both sides

OpenStudy (anonymous):

Okay -4x=-12. Now do I multiply each side by -1 now and the answer would be x=3?

OpenStudy (anonymous):

Yes but you don't have to do the -1 stuff just divide both by -4 to get\[x=3\]

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

Oh thats i forgot thanks :)

OpenStudy (anonymous):

np